OpenStudy (anonymous):

A child is creating a pyramid with building blocks. The top three levels include 4 blocks, 10 blocks, and 16 blocks. How many blocks would be needed for a pyramid 25 levels tall? Use complete sentences to explain how a sum of an arithmetic series was applied? Thanks!!

OpenStudy (anonymous):

PEDMAS s(25) = a/2 [2a + ( - 1)d] 25/2 [2 * 4 + (25 - 1)*6] 25/2 [8 + 25 * 6] 25/2 * 152 = 1,900 Hope this helps :D

OpenStudy (anonymous):

Thanks, do you know a simple equation I could give.. for example something like 4+6n where n= the level??